40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Feb 5 21:42:28 2009, Yoichi, Update, PSL, My thoughts on ISS INMONPD_Spectrum_1-10Hz.pdf
    Reply  Fri Feb 6 09:52:35 2009, Kakeru, Update, PSL, Current shunt transfar function TF_CS_gain.pngTF_CS_phase.png
       Reply  Fri Feb 6 23:23:48 2009, Kakeru, Yoichi, Update, PSL, ISS is fixed Current_ShuntTF_gain.pngCurrent_ShuntTF_phase.png
          Reply  Mon Feb 9 16:02:42 2009, Yoichi, Update, PSL, PSL relative intensity noise RIN-13dB.pngRIN-21dB.png
             Reply  Tue Feb 10 23:36:25 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
                Reply  Wed Feb 11 07:28:25 2009, Yoichi, Update, PSL, PA current and laser output 
                   Reply  Wed Feb 11 23:51:53 2009, Kakeru, Update, PSL, PA current and laser output PA_current_output.png
Message ID: 1291     Entry time: Wed Feb 11 07:28:25 2009     In reply to: 1289     Reply to this: 1295
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: PA current and laser output 
I think we should also plot the laser power at the MOPA output. The horizontal axis should be the absolute current value read from the PA current monitor channel, not the slider value.

This result is consistent with my hypothesis that the thermal effect is canceling the power change at low frequencies (see elog:1276).
But if it is really caused by thermal effect or not is still unknown.

I'd like to see a larger scan into the lower current region.

I changed the PA current and measured laser output power (monitor PD signal).
The gain of ISS is 13dB
Attached figure is the relation of PA current and the average and standard diviation of laser output.
The average of output power decreas as current increase. It looks something is wrong with PA.
When current is -0.125, 0, 0.5, ISS become ocsilating. This looks to be changed from previous measurement.

I wrote matlab code for this measurement. The code is
This function uses
ELOG V3.1.3-